OpenStudy (anonymous):

What is the restriction on the product of x^2 + 3x = 2 / x^2 - 2x - 3 multiplied by x^2 + 4x + 3 / x + 2

OpenStudy (fibonaccichick666):

that equals sign in there doesn't make much sense, can you double check and rewrite?

OpenStudy (danjs):

use parenthesis

OpenStudy (anonymous):

Sorry, it's \[ \frac{ x^2 + 3x + 2 }{ x^2 - 2x - 3 } \times \frac{ x^2 + 4x + 3 }{ x + 2 }\]

OpenStudy (fibonaccichick666):

okk better. So, can you simplify at all to start?

OpenStudy (anonymous):

The numerator of the first fraction can be simplified into (x + 2)(x + 1), the denominator can be simplified into (x - 3)(x + 1), the numerator of the second fraction can be simplified into (x + 3)(x + 1)

OpenStudy (fibonaccichick666):

good! so you can cancel like polynomials out, what does that leave you with?

OpenStudy (anonymous):

It leaves me with the (x + 2), (x - 3), (x + 3), and (x + 1) right?

Directrix (directrix):

I'm thinking that the restrictions on the product would stem from values of x that cause either denominator of the original expression to be 0. (Division by zero is undefined.) The first denominator factors to: ( x + 1) and (x - 3) The second denominator is just: (x + 2) To get the restricted values of x, solve each of the following for x: (x + 1) = 0 (x - 3) = 0 (x + 2) = 0 @maciforster

OpenStudy (fibonaccichick666):

@Directrix , restriction usually means make it so we dont have an indeterminant or undefined form

OpenStudy (fibonaccichick666):

and it should still leave you with an equation. Not a listing of binomials

OpenStudy (danjs):

\[\frac{ (x+2)(x+1) }{ (x-3)(x+1) } * \frac{ (x+3)(x+1) }{ (x+2) } = \frac{ (x+2)(x+1)(x+3)(x+1) }{ (x-3)(x+1)(x+2) }\] As you can see in the right hand side of above, you are able to cancel out (x+2) and (x+1) from both numerator and denominator. When you cancel a quantity from a fraction or equation, you need to keep track of the RESTRICTION on the variable. In this case both the quantities (x+2) and (x+1) can not be zero. Sometimes when you solve equations, quantities cancel out, and you need to keep track of that in your final answer.

OpenStudy (danjs):

For instance, if you have: \[\frac{ (6x+6)(x+3) }{ (x+3) } > 35\] After canceling out the (x+3), you are left with: \[6x + 6 > 35\] -AND- \[(x+3) \neq 0\] The second is your restriction. So the final answer would be: \[x > \frac{ 29 }{ 6 } ,and; x \neq -3\]

OpenStudy (danjs):

in that case the restriction is not part of the solution, so it is not needed. If x<29/6 then you would need to include the \[x \neq -3\] in your solution as a restriction.

OpenStudy (anonymous):

Thank you!!!

Directrix (directrix):

@FibonacciChick666 Read the sentence above the "listing of binomials." To get the restricted values of x, solve each of the following for x: (x + 1) = 0 (x - 3) = 0 (x + 2) = 0 The set of linear equations is for the Asker to solve to find the restricted values of x. x = -1, x = 3, and x = -2 are restricted values.

OpenStudy (fibonaccichick666):

I'm sorry, I read numerator like 12 times